OpenStudy (bananas1234):

Identify the restrictions on the domain. x+2/x-5 divided by x-6/x

OpenStudy (anonymous):

First, didvide the two expressions: \[\Huge \frac{ \frac{ x+2 }{ x-5 } }{ \frac{ x-6 }{ x} }\]

OpenStudy (bananas1234):

What do i do next?

OpenStudy (anonymous):

I'm trying to load it, it doesnt seem to be working

OpenStudy (bananas1234):

ok

OpenStudy (anonymous):

\[\Large \frac{x+2}{x-5} \times \frac{x}{x-6} \]

OpenStudy (anonymous):

You have to multiply by the reciprocal of x-6/x when you divide, btw. Now you have a new expression \[\frac{(x+2)x}{(x-5)(x-6)}\]

OpenStudy (bananas1234):

i see

OpenStudy (anonymous):

The restrictions of the domain of the function basically mean what number we cannot plug into our equation (what makes the denominator 0 or what makes the numerator unrea, so if there was a sqrt up there, what makes a negative number, but here isn't one so we don;t have to worry about the numerator). Setting the denominator equal to 0 we get (x-5)(x-6)=0 x=5 x=6 These are your restrcitions

OpenStudy (freckles):

x=0 would also need to be excluded from the domain

OpenStudy (bananas1234):

so what does x not equal?

OpenStudy (freckles):

\[f(x)=\frac{x+2}{x-5} \div \frac{x-6}{x}\] like for example if you plug in 0 \[f(0)=\frac{0+2}{0-5} \div \frac{0-6}{0}\] what would you say that output is

OpenStudy (bananas1234):

0?

OpenStudy (freckles):

division by 0 does not give you 0

OpenStudy (bananas1234):

6?

OpenStudy (freckles):

division by 0 is not good it is undefined a/0 is undefined

OpenStudy (freckles):

anyways I hope you can see what numbers to exclude from the domain

OpenStudy (bananas1234):

so x is not 5 and 6?

OpenStudy (freckles):

or also?

OpenStudy (freckles):

there are 3 values we talked about

OpenStudy (freckles):

2 given by the first guy and then I gave you another remember we cannot divide by 0

OpenStudy (freckles):

\[f(x)=\frac{x+2}{x-5} \div \frac{x-6}{x} \] do you see an x on the bottom of the second fraction?

OpenStudy (freckles):

that cannot be 0 because we cannot divide by 0

OpenStudy (bananas1234):

my options are x is not 5,6 x is not 5,0 x is not -2,6 x is not -2, 0

OpenStudy (freckles):

well I would go with option e then because x cannot be 0,5,6

OpenStudy (bananas1234):

a?

OpenStudy (freckles):

no e

OpenStudy (freckles):

none of the above

OpenStudy (freckles):

x is cannot be 0,5,6

OpenStudy (bananas1234):

hmmm im sorry but there is only a,b,c, and d

OpenStudy (freckles):

well I would talk to your teacher because f(0) and f(5) and f(6) are all not defined

OpenStudy (bananas1234):

ok

OpenStudy (freckles):

\[f(x)=\frac{x+2}{x-5} \div \frac{x-6}{x} \\f(0)=\frac{0+2}{0-5} \div \frac{0-6}{0}=\frac{2}{-5} \div \frac{-6}{0} \text{(second fraction undefined } \\ f(5)=\frac{5+2}{5-5} \div \frac{5-6}{5} =\frac{7}{0} \div \frac{-1}{5} \text{ (first fraction undefined } \\ f(6)=\frac{6+2}{6-5} \div \frac{6-6}{6}=\frac{8}{1} \div \frac{0}{6}=8 \div 0=\frac{8}{0} \text{(again we cannot divide by 0}\] if your teacher doesn't know write all of this out and tell you cannot divide by 0

OpenStudy (bananas1234):

ok, thank you for the info. what option would you go with. none of them seem to be completely correct, but i have to chose one, and then i will talk to my teacher about it later. :)

OpenStudy (freckles):

i guess the 5,6 one I don't know

OpenStudy (bananas1234):

Ok, i will talk to her about it, thanks

OpenStudy (freckles):

hey and the expression was really what I wrote ,right?

OpenStudy (freckles):